### v6

# Less is More: Token-Efficient Video-QA via Adaptive Frame-Pruning and Semantic Graph Integration

Abstract
The practical application of Multimodal Large Language Models (MLLMs) to Video Question Answering (Video-QA) is severely hindered by the high token cost of processing numerous video frames. While increasing the number of sampled frames is a common strategy, we observe a ”less is more” phenomenon where excessive frames can paradoxically degrade performance due to context dilution. Concurrently, state-of-the-art keyframe selection methods, while effective, still yield significant temporal redundancy, which we term ’visual echoes’. To address these dual challenges, we propose Adaptive Frame-Pruning (AFP), a novel post-processing method that intelligently prunes the selected keyframes. AFP employs an adaptive hierarchical clustering algorithm on a fused ResNet-50 and CLIP feature space to identify and merge these echoes into single representatives. To compensate for information loss, we then introduce a lightweight, text-based semantic graph that provides critical context with minimal token overhead. Conducting extensive experiments on the LongVideoBench and VideoMME benchmarks across multiple leading MLLMs, our full approach demonstrates a drastic reduction in required frames by up to 86.9% and total input tokens by up to 83.2%. Crucially, by providing a concise, high-quality set of frames, our method not only enhances efficiency but often improves accuracy over baselines that use more frames. The code will be released upon publication.
Abstract
This supplementary material offers extensive validation and ensures the full reproducibility of our main paper on Adaptive Frame-Pruning (AFP). We begin by presenting detailed results from our comprehensive ablation studies—covering components, prompt formulation, and hyperparameter sensitivity—which collectively justify our final model design. To demonstrate the robustness and versatility of our approach, we include a generalizability study on the T* selector and provide additional qualitative examples for different QA tasks. Furthermore, we offer in-depth implementation details, from algorithmic parameters to environment setup, to facilitate replication. The document concludes with a transparent discussion of our method’s limitations and a failure case analysis, outlining clear directions for future research.
Introduction
Figure 1: The ”Less is More” phenomenon is a recurring challenge in Video Understanding. The plot aggregates findings from our own experiments (on LongVideoBench and VideoMME) and recent literature (DrVideo on EgoSchema). Across different models and datasets, performance consistently peaks at a small number of frames and then degrades, highlighting a universal need for intelligent frame pruning over simple accumulation.
The advent of Multimodal Large Language Models (MLLMs), such as GPT-4o (Hurst et al. 2024), Gemini (Team et al. 2023), Video-Chatgpt (Maaz et al. 2023), MovieChat (Song et al. 2024b) and Video-llava (Lin et al. 2023) has catalyzed a paradigm shift in vision-language understanding. These models demonstrate remarkable capabilities in performing complex reasoning over visual and textual inputs, with Video Question Answering (Video-QA) emerging as a key application. However, a significant bottleneck impedes their widespread adoption: the prohibitive token cost associated with processing long video sequences.
A common strategy to mitigate this is to sample a subset of frames. However, a counter-intuitive phenomenon emerges: simply increasing the number of used frames does not guarantee better performance. As illustrated in Figure 1, multiple experiments in state-of-the-art keyframe search strategies such as T* (Ye et al. 2025) and VSLS (Guo et al. 2025), performance that the accuracy of Video-QA often peaks at a relatively small number of frames (e.g., 8) and then degrades as more frames are added. Crucially, this ”less is more” phenomenon is not an isolated finding. Recent state-of-the-art work, such as DrVideo (Ma et al. 2025) has independently reported the exact same trend, where increasing the number of top-k keyframes from 5 to 20 led to a consistent decline in QA accuracy on the EgoSchema benchmark. This recurring observation across different models, datasets, and keyframe selection methods points to a fundamental challenge in the field, which we term ”visual echoes”: even advanced keyframe selection methods often yield a set of frames with significant internal redundancy, which may introduce distracting noise and overwhelm the MLLM’s reasoning capacity. This raises a pivotal research question: For a Video-QA instance, what is the optimal, concise set of frames for a given query?
To address this specific challenge, we propose a novel, two-pronged approach designed to drastically enhance the efficiency of Video-QA. First, we introduce Adaptive Frame-Pruning (AFP), a versatile post-processing method designed to serve as a refinement layer on top of any keyframe selector. AFP employs an adaptive hierarchical clustering algorithm, which operates on a fused feature space of ResNet (He et al. 2016)and CLIP (Radford et al. 2021) embeddings, to intelligently identify and consolidate ‘visual echoes’ into single, representative frames. Recognizing that such an aggressive reduction might lead to the loss of subtle but crucial details, we introduce our second component: a lightweight, text-based semantic graph. This graph explicitly outlines the key objects and their relationships relevant to the QA context, compensating for potential information loss. Crucially, this semantic context can often be repurposed from the intermediate outputs of advanced selectors, incurring negligible additional cost. As illustrated in Figure 2, our approach transforms a high-cost, redundant initial prompt into a highly efficient, optimized one.
The main contributions of our work are threefold:
  • We empirically demonstrate and analyze the ”less is more” phenomenon in Video-QA, providing a strong motivation for frame pruning over simple frame selection.
  • We propose a novel algorithm, Adaptive Frame-Pruning (AFP), to effectively eliminate the ‘visual echoes’ (redundancy) present in the output of state-of-the-art keyframe selectors.
  • We conduct extensive experiments on the LongVideoBench and VideoMME benchmarks across a diverse set of MLLMs, demonstrating that our method not only enhances efficiency but often improves accuracy by mitigating context dilution.
Figure 2: Conceptual Overview. (a) An initial prompt with redundant ‘visual echoes’ has a high token cost. (b) Our AFP algorithm prunes these frames and combines them with a semantic graph to form an optimized, low-cost prompt.
Related Work
Keyframe Selection for Efficient Video-QA. The high token cost of applying Multimodal Large Language Models (MLLMs) (Yan et al. 2025; Hurst et al. 2024) to long videos has made keyframe selection a crucial preprocessing step. Approaches range from learnable, end-to-end retrieval models like R-VLM (Xu et al. 2023) to more flexible training-free methods. The latter, often formulating selection as a query-aware optimization problem (Tang et al. 2025; Wang et al. 2025; Tang et al. 2025; Liang et al. 2024), have seen significant advances. Seminal works like T* (Ye et al. 2025) and VSLS (Guo et al. 2025) perform sophisticated temporal searches by leveraging spatial detectors and semantic-logical relations. However, a key challenge emerges from these advanced selectors: in their effort to be comprehensive, they often produce a keyframe set with significant internal redundancy—the ‘visual echoes’ we aim to address. Our Adaptive Frame-Pruning (AFP) is positioned not as a replacement for these selectors, but as a crucial subsequent refinement stage that prunes this redundancy.
Redundancy and Representation in MLLM Input. The ”less is more” principle has become a focal point in MLLM efficiency research. In the image domain, methods like TRIM (Song et al. 2024a) perform token-level reduction by selecting relevant image patches. While sharing this philosophy, our AFP operates at the frame level, addressing the unique challenge of temporal redundancy in videos. Instead of patch-level selection, we employ adaptive clustering on the entire frame’s features to consolidate ‘visual echoes’ arising from continuous motion. By operating at the frame level, our approach is tailored to the specific nature of video data, tackling the challenge of redundancy directly, a dimension not addressed by patch-level image pruning methods.
Semantic Compensation. To compensate for information loss from pruning, we leverage semantic representations. Traditionally, structured information like scene graphs (Krishna et al. 2017) requires complex Graph Neural Network (GNN) (Wu et al. 2020) architectures. A key innovation of our work is to circumvent this by proposing a textualized semantic graph. This allows us to inject rich, structured context directly into the MLLM’s prompt with minimal token overhead, a novel and highly compatible approach. Crucially, as detailed in our methodology, this graph can often be generated with zero additional cost by repurposing intermediate artifacts from advanced selectors like VSLS.
Methodology
Our approach enhances Video-QA efficiency via a two-pronged strategy of Pruning and Compensation. We first prune the visual stream to a concise, non-redundant keyframe set, then compensate for information loss by injecting an efficient, abstract semantic layer. The two main stages, illustrated in Figure 3, are: (1) Adaptive Frame-Pruning (AFP) (pruning) and (2) Textual Semantic Graph Integration (compensation).
Figure 3: The Overall Pipeline of Our Proposed Method. An upstream selector provides initial frames. Our Adaptive Frame-Pruning (AFP) module then takes over, performing (1) fused feature extraction and adaptive clustering to produce representative keyframes, and (2) concurrent semantic graph generation. Both are combined into an optimized prompt for the MLLM.
Adaptive Frame-Pruning (AFP)
The core of our frame reduction strategy is AFP, an algorithm designed to identify and consolidate ‘visual echoes’. The process involves three key steps: extracting robust fused features, performing adaptive clustering, and selecting a representative frame from each cluster. The entire process is formalized in Algorithm 1.
Fused Feature Extraction.
To ensure that our clustering is sensitive to both low-level visual patterns and high-level semantic content, we employ a fused feature representation. For each keyframe, we extract features from two powerful, pre-trained models: a ResNet-50 (He et al. 2016) backbone, renowned for its strong performance on visual recognition tasks, and a CLIP ViT-B/32 model (Radford et al. 2021), which excels at aligning images with text and capturing rich semantic meaning. The high-dimensional outputs from both models are passed through separate linear projection layers to map them to a shared 512-dimensional space. The final feature vector $\mathbf{f}_{\text{fused}}$ for a frame is a weighted combination of the L2-normalized projected features:
$\mathbf{f}_{\text{fused}}=(1-\alpha)\cdot\mathbf{f}_{\text{ResNet}}+\alpha\cdot\mathbf{f}_{\text{CLIP}},$ (1)
where $\alpha$ is a fusion ratio hyperparameter, set to 0.6 in our experiments to slightly favor the semantic features from CLIP.
Adaptive Hierarchical Clustering.
A key innovation of AFP is its ability to adapt the clustering process to the specific content of each video, avoiding a fixed, pre-defined number of clusters. We use Agglomerative Hierarchical Clustering (Murtagh and Contreras 2012), which requires a distance metric and a linkage criterion.
Distance Metric. Our distance metric jointly considers visual similarity and temporal proximity. For any two frames $i$ and $j$ with fused features $\mathbf{f}_{i},\mathbf{f}_{j}$ and timestamps $t_{i},t_{j}$, the combined distance $D(i,j)$ is defined as:
$D(i,j)=\beta\cdot d_{\text{cos}}(\mathbf{f}_{i},\mathbf{f}_{j})+(1-\beta)\cdot d_{\text{temp}}(t_{i},t_{j}),$ (2)
where $d_{\text{cos}}$ is the cosine distance, $d_{\text{temp}}$ is the normalized absolute difference of timestamps, and $\beta$ is a weighting factor (set to 0.7). This combined metric ensures that frames that are visually similar but temporally distant are less likely to be clustered together.
Adaptive Distance Threshold. Instead of setting a fixed number of clusters, we use a dynamic distance threshold $\tau$. This threshold is determined by applying a Gaussian Kernel Density Estimator (KDE) to the distribution of all pairwise visual distances ($d_{\text{cos}}$) within the keyframe set. We identify the peak of this density distribution, which corresponds to the most common distance between frames, and set $\tau$ slightly above this value. This allows the algorithm to automatically find a natural grouping of frames based on their inherent similarity structure for that specific video.
Representative Frame Selection.
After clustering, a single representative frame must be selected from each resulting cluster. This is a critical step, as different selection strategies can impact performance. We conducted a detailed ablation study on several strategies (see Appendix for full details). Our findings show that a centroid-based selection approach is the most robust and effective. Therefore, for all our main experiments, we identify the ”centroid frame” for each cluster—the one with the minimum average visual feature distance to all other frames within that same cluster. This self-contained strategy selects the most visually representative frame, minimizing information loss during pruning without relying on external scores and thus enhancing the generalizability of our AFP framework.
Algorithm 1 Adaptive Frame-Pruning (AFP)
Input: Initial keyframes $K=\{k_{1},...,k_{N}\}$, relevance scores $S=\{s_{1},...,s_{N}\}$, timestamps $T=\{t_{1},...,t_{N}\}$.
Parameters: Fusion ratio $\alpha$, distance weight $\beta$.
Output: Refined keyframes $K^{\prime}$.
1: Initialize feature set $F\leftarrow\emptyset$.
2:for each keyframe $k_{i}\in K$ do
3:  Extract and project ResNet and CLIP features.
4:  Compute $\mathbf{f}_{\text{fused},i}$ using Eq. 1.
5:  $F\leftarrow F\cup\{\mathbf{f}_{\text{fused},i}\}$.
6:end for
7: Compute pairwise combined distance matrix $D$ using Eq. 2.
8: Compute pairwise visual distances $D_{\text{cos}}$.
9: Apply KDE to $D_{\text{cos}}$ to find adaptive threshold $\tau$.
10: Perform Agglomerative Clustering on $K$ using distance matrix $D$ and threshold $\tau$ to get clusters $C=\{C_{1},...,C_{M}\}$.
11: Initialize refined set $K^{\prime}\leftarrow\emptyset$.
12:for each cluster $C_{j}\in C$ do
13:  Find $k^{*}=\arg\max_{k_{i}\in C_{j}}s_{i}$.
14:  $K^{\prime}\leftarrow K^{\prime}\cup\{k^{*}\}$.
15:end for
16:return $K^{\prime}$
Textual Semantic Graph Integration
To counteract potential information loss from the aggressive frame pruning performed by AFP, we introduce a supplementary, low-cost semantic context in the form of a textualized semantic graph. A key novelty of our work is the flexible and highly efficient manner in which this graph is generated, for which our framework supports a dual-strategy to ensure both maximum efficiency and universal applicability.
Case 1: Repurposing from Advanced Selectors (Zero-Cost). Our primary approach, used in all experiments involving VSLS and T*, is to capitalize on its intermediate artifacts. Methods like VSLS often deconstruct a query into key objects and logical relations to guide their search. We simply intercept and repurpose this valuable, untapped resource. By textualizing these pre-existing relations, we generate the semantic graph with zero additional extraction cost, as illustrated by the template in Figure 4.
Case 2: LLM-based Generation for General Applicability (Low-Cost). To ensure our framework’s universal applicability to any keyframe selector (even those that do not provide semantic relations, like AKS (Tang et al. 2025) or Uniform Sampling), we also propose a lightweight, fallback generation strategy. In this scenario, we prompt a large model (e.g., GPT-4o) with only the textual content of the Video-QA task—the question and options. Unlike previous methods that restrict relation types, our carefully designed prompt (detailed in the Appendix) instructs the MLLM to act as an ”expert visual scene analyst” and infer a wide range of plausible, open-ended relationships. This process is extremely token-efficient, as it involves no image inputs, consuming a negligible amount of tokens.
As our ablation studies demonstrate (see Section 5), the performance boost from integrating the semantic graph is substantial. This dual-strategy ensures that our AFP + Graph framework is not only powerful but also highly versatile, offering a significant performance gain for a minimal (and often zero) computational investment.
Figure 4: The template of our textualized semantic graph. This concise text block, which can be generated either by repurposing or via a low-cost LLM call, is inserted directly into the downstream Video-QA period as a part of the prompt for MLLM to provide high-level semantic context.
Experiments
Experimental Setup
Datasets and Models. We evaluate our method on two challenging long-video question-answering benchmarks: LongVideoBench (Ye, Wang, and Sun 2025) and VideoMME (Fu et al. 2024). To demonstrate the generalizability of our approach, we conduct experiments across three distinct MLLMs: the proprietary GPT-4o (Hurst et al. 2024), and two powerful open-source models, Qwen2.5-VL-7B-Instruct (Team 2025) and LLaVA-Video-7B-Qwen2 (Li et al. 2024).
Compared Methods and Baselines. Our evaluation protocol is designed to rigorously assess our method. For each MLLM, we compare three keyframe sampling strategies in our main results (Table 1):
  • Uniform Sampling: A naive baseline where a fixed number of frames (32 or 8) are sampled at regular intervals.
  • VSLS: A sophisticated query-aware keyframe selector. It serves as our primary upstream selector and a strong baseline. VSLS deconstructs the query into key entities and semantic-logical relations, using them to score and select the most relevant frames.
  • AFP + Graph: Our full method, which takes the output of the upstream selector as input and applies our adaptive pruning and semantic graph compensation.
In this paper, we primarily use VSLS as the upstream selector to demonstrate our method’s effectiveness. To further validate the generalizability of AFP, we also provide results of applying it to another state-of-the-art keyframe selector, T* (Ye et al. 2025), in the appendix.
Evaluation Metrics. We evaluate all methods based on two primary criteria:
  • Effectiveness: We use the official Accuracy (%) metric for both benchmarks, calculated via strict exact matching of the predicted option letter against the ground truth.
  • Efficiency: We measure the Average Frames used per query as the metric for efficiency.
A detailed token analysis for GPT-4o is also provided.
Implementation Details and Hyperparameter Analysis. To ensure reproducibility, we detail our implementation. Our method’s core hyperparameters, the feature fusion ratio $\alpha$ and distance weight $\beta$, were determined through a systematic sensitivity analysis, the results of which are visualized in Figure 5. This analysis reveals a clear cost-utility trade-off, allowing users to tune the parameters to prioritize either performance or efficiency. For our main experiments, we selected $\alpha=0.6$ and $\beta=0.9$, as this configuration provides the best balance of high accuracy and high efficiency across our development experiments. A detailed table and analysis are provided in Appendix D. In addition, the detailed MLLM prompt structure is detailed in Appendix B.
Figure 5: Cost-Utility trade-off analysis for hyperparameters $\alpha$ and $\beta$. Cost (left y-axis, red) is measured by the average number of output frames, while Utility (right y-axis, blue) is measured by the average QA accuracy on a subset of LongVideoBench. The plots demonstrate that our parameters are tunable knobs: users can adjust them to prioritize either higher accuracy (e.g., lower $\beta$) or maximum efficiency (e.g., higher $\beta$). Our default choice ($\alpha=0.6,\beta=0.9$‘) represents a strong balance.
Main Results and Analysis
Our analysis begins by investigating the ”less is more” phenomenon inherent in MLLM-based Video-QA. We then present our main results, demonstrating how our proposed method effectively navigates the critical trade-off between performance and efficiency.
The ”Less is More” Paradox in Baselines. Before analyzing our method’s performance, it is crucial to first observe the ”less is more” phenomenon. In addition to the trend visualized in Figure 1, our main results in Table 1 further corroborate this observation, which serves as a key motivation for our work. A common assumption is that more frames lead to better performance, but this is not always true. For example, with GPT-4o + VSLS on VideoMME, increasing the input from 8 to 32 frames results in a performance drop on long videos (from 56.9% to 55.2%). Similarly, with Qwen2.5-VL-7B-Instruct + VSLS on LongVideoBench, increasing from 8 to 32 frames causes a significant drop in long videos (from 45.8% to 38.7%). This ”context dilution” paradox, where excessive frames introduce noise, underscores the need not just for relevant frames, but for a concisely pruned, optimal set.
LongVideoBench VideoMME Model and Method Avg. Video Length Accuracy (%) Model and Method Avg. Video Length Accuracy (%) Frames Long Medium Short Frames Long Medium Short (900-3600s) (180-600s) (15-60s) (30-60min) (4-15min) (0-2min) Evaluation starting from 32 Keyframes GPT-4o + Uniform 32.0 53.8 56.5 74.0 GPT-4o + Uniform 32.0 55.2 61.0 71.4 GPT-4o + VSLS 32.0 54.2 60.0 76.0 GPT-4o + VSLS 32.0 55.2 61.9 71.9 GPT-4o + AFP+Graph 4.2 49.4 51.5 80.0 GPT-4o + AFP+Graph 4.3 53.0 56.4 63.4 Qwen2.5-VL-7B + Uniform 32.0 32.7 36.5 50.0 Qwen2.5-VL-7B + Uniform 32.0 37.5 39.9 54.1 Qwen2.5-VL-7B + VSLS 32.0 38.7 42.3 54.0 Qwen2.5-VL-7B + VSLS 32.0 37.9 50.0 55.8 Qwen2.5-VL-7B + AFP+Graph 4.2 42.9 46.9 66.0 Qwen2.5-VL-7B + AFP+Graph 4.3 39.9 44.0 51.8 LLaVA-Video-7B + Uniform 32.0 42.3 45.8 54.0 LLaVA-Video-7B + Uniform 32.0 35.9 36.5 37.4 LLaVA-Video-7B + VSLS 32.0 41.7 48.1 54.0 LLaVA-Video-7B + VSLS 32.0 36.9 39.0 39.5 LLaVA-Video-7B + AFP+Graph 4.2 45.2 50.0 62.0 LLaVA-Video-7B + AFP+Graph 4.3 43.8 47.9 56.7 Evaluation starting from 8 Keyframes GPT-4o + Uniform 8.0 47.1 49.4 67.3 GPT-4o + Uniform 8.0 55.2 60.2 69.6 GPT-4o + VSLS 8.0 51.2 58.9 74.0 GPT-4o + VSLS 8.0 56.9 60.7 68.2 GPT-4o + AFP+Graph 2.2 47.3 53.5 84.0 GPT-4o + AFP+Graph 2.1 53.5 56.5 63.8 Qwen2.5-VL-7B + Uniform 8.0 41.1 43.1 62.0 Qwen2.5-VL-7B + Uniform 8.0 38.0 47.3 55.4 Qwen2.5-VL-7B + VSLS 8.0 45.8 49.2 54.0 Qwen2.5-VL-7B + VSLS 8.0 43.2 49.6 60.8 Qwen2.5-VL-7B + AFP+Graph 2.2 42.3 44.2 64.0 Qwen2.5-VL-7B + AFP+Graph 2.1 38.1 40.2 50.1 LLaVA-Video-7B + Uniform 8.0 42.0 46.5 50.0 LLaVA-Video-7B + Uniform 8.0 38.0 39.7 38.2 LLaVA-Video-7B + VSLS 8.0 42.3 46.9 50.0 LLaVA-Video-7B + VSLS 8.0 38.5 38.5 38.5 LLaVA-Video-7B + AFP+Graph 2.2 45.5 49.6 72.0 LLaVA-Video-7B + AFP+Graph 2.1 44.2 47.9 54.5
Table 1: Main Results on LongVideoBench and VideoMME with different MLLMs. We compare our full method (AFP + Graph) against a sophisticated baseline (VSLS) and a naive baseline (Uniform Sampling). Our method achieves superior or highly competitive accuracy while using drastically fewer frames (see Avg. Frames col.). Bold indicates the best performing method within each 3-row comparison block for each accuracy column.
Main Quantitative and Visual Analysis. Our main experimental results are presented in Table 1. This comprehensive table compares our full method (‘AFP + Graph ‘) against two baselines (‘Uniform Sampling‘ and ‘VSLS ‘) across three different MLLMs. A key aspect of this table is the ‘Avg. Frames‘ column: for the ‘Uniform‘ and ‘VSLS ‘ baselines, this value is a fixed integer (e.g., 32.0 or 8.0) as every instance in the test set uses that exact number of frames. In contrast, for our ‘AFP + Graph ‘ method, this column reports a decimal value, representing the average number of frames output by our adaptive algorithm across all instances, which varies depending on the video’s content. The subheadings (e.g., ”Evaluation starting from 32 Keyframes”) define the fixed frame count for the baselines and the input frame count for our method within each block. The subheadings (e.g., ”Evaluation starting from 32 Keyframes”) define the fixed frame count for the ‘Uniform‘ and ‘VSLS ‘ baselines, and simultaneously specify the input frame count for our ‘AFP + Graph ‘ method.
AFP’s Superiority Across Diverse MLLMs. Having established the challenge of context dilution, we now analyze the superiority of our AFP + Graph method. The results unequivocally demonstrate that our approach consistently delivers the best performance among all methods, especially when considering the efficiency-accuracy trade-off. Our method achieves this by combining intelligent pruning with a semantic graph that, as detailed in our Methodology, is repurposed from the upstream selector with no extra computational overhead.
The most compelling evidence of our method’s value comes from its performance with these open-source MLLMs. For instance, using LLaVA-Video-7B-Qwen2 on LongVideoBench (starting from 8 frames), our method uses only an average of 2.2 frames yet boosts the accuracy on short videos from 50.0% (VSLS) to a remarkable 72.0%—a 22-point absolute improvement. This massive gain in both effectiveness and efficiency is a consistent trend across both Qwen2.5-VL-7B-Instruct and LLaVA-Video-7B-Qwen2. This suggests that open-source models are particularly susceptible to context dilution, and our AFP + Graph method is highly effective at providing them with a cleaner, more focused visual and semantic context, thereby unlocking their true potential.
Figure 6: Average token consumption comparison across datasets and methods. Our method (AFP + Graph) consistently and drastically reduces the token requirements compared to the VSLS baseline across all settings and on both LongVideoBench and VideoMME, highlighting its superior and generalizable efficiency.
Even with the more robust GPT-4o, our method provides a highly favorable efficiency-performance trade-off. While VSLS often sets a high performance bar, our method achieves highly competitive results with a fraction of the input. This immense efficiency gain is best visualized in Figure 6, which details the token consumption for GPT-4o. The chart clearly shows that our ‘AFP + Graph ‘ approach reduces the token cost by a significant margin in all settings. For example, on LongVideoBench (from Top-32), the token count plummets from 2980 with VSLS to just 609 with our method, a reduction of nearly 5x.
Figure 7: Efficiency and Performance results of AFP + Graph on VSLS Keyframes across different MLLMs. The plot shows the trade-off between overall accuracy and the average number of frames used on both LongVideoBench (left) and VideoMME (right) datasets. Solid lines represent the high-cost VSLS baseline at 8 and 32 frames. Dashed lines represent our low-cost AFP + Graph method applied to the corresponding VSLS outputs. Our method consistently operates in the top-left (high-accuracy, low-frame-count) region, demonstrating its effectiveness as a universal refinement module.
Key Visualization of the main results. Other than main results presented in Table 1, the key findings are also visualized in the performance-efficiency plot in Figure 7. The ideal method should reside in the top-left corner of the plot, signifying high accuracy with low frame usage. Our analysis across diverse MLLMs reveals two major strengths of our AFP + Graph approach.
1. Unlocking Performance for Open-Source Models. The most compelling evidence of our method’s value is its profound impact on open-source models. As clearly shown in Figure 7, for both LLaVA-Video-7B-Qwen2 and Qwen2.5-VL-7B-Instruct, our method’s data points (dashed lines) are consistently located not only far to the left (more efficient) but also significantly above (more effective) the VSLS baseline points (solid lines). For instance, on LongVideoBench, our method boosts LLaVA’s accuracy from 48.1% (at 8 frames) to 51.5% (at only 2.5 frames). This demonstrates that our intelligent pruning does more than save costs—it actively mitigates the ”context dilution” problem, to which these models appear more susceptible, thereby unlocking their true potential by providing a cleaner, more focused visual context.
2. Achieving Favorable Trade-offs for State-of-the-Art Models. For the highly robust GPT-4o, while the VSLS baseline with a full frame set often establishes the performance ceiling, our AFP + Graph method offers an extremely favorable efficiency-performance trade-off. Figure 7 shows our method’s points positioned far to the left, achieving accuracies that are highly competitive with the baseline while using only a fraction of the frames (e.g., achieving 52.1% vs. baseline’s 58.1% on LongVideoBench with 2.5 vs. 8 frames). This immense efficiency gain is further highlighted by the token analysis in Figure 6. The plot clearly shows that our approach slashes the token cost by a factor of 4-5x in high-frame-count scenarios, such as reducing the input for LongVideoBench (from Top-32) from 2,980 to just 609 tokens. By drastically reducing costs for a moderate and often acceptable performance trade-off, our method makes large-scale Video-QA applications with expensive proprietary models economically viable.
Dataset Model Method Avg. Frames Long (%) Medium (%) Short (%) Ablation from VSLS Top-32 Keyframes VideoMME GPT-4o Uniform (Matched) 4.3 52.3 52.6 57.1 VSLS (Top-N, Matched) 4.3 49.7 50.2 54.1 AFP only (Ours) 4.3 52.5 52.5 57.7 AFP + Graph (Ours) 4.3 53.0 56.4 63.4 VideoMME Qwen2.5-VL-7B Uniform (Matched) 4.3 39.1 43.5 43.6 VSLS (Top-N, Matched) 4.3 37.9 40.2 42.5 AFP only (Ours) 4.3 38.3 39.8 42.5 AFP + Graph (Ours) 4.3 39.9 44.0 51.8 Ablation from VSLS Top-8 Keyframes LongVideoBench LLaVA-Video-7B Uniform (Matched) 2.2 39.3 46.2 42.0 VSLS (Top-N, Matched) 2.2 40.5 42.7 46.0 AFP only (Ours) 2.2 41.1 44.6 44.0 AFP + Graph (Ours) 2.2 45.5 49.6 72.0
Table 2: Component-wise ablation study across different MLLMs and datasets under a matched frame budget.AFP + Graph ‘ consistently outperforms all other low-budget strategies, highlighting the synergy between intelligent clustering and semantic compensation.
Experiment Analysis and Ablation Studies
Having established the overall superiority of our method in the main experiments, we now conduct a more focused analysis to dissect the individual contributions of our core components: the clustering-based selection (AFP) and the semantic graph compensation.
Component-wise Ablation
To isolate the effect of each component, we present a comprehensive ablation study in Table 2. This study evaluates four distinct low-frame-budget strategies across a variety of models, datasets, and initial keyframe settings.
Ablation Strategies. All four strategies are designed to operate under a strictly matched frame budget for each individual Video-QA instance. For each setting, the ‘Avg. Frames‘ column in Table 2 reports the average number of frames used per query across the entire evaluation subset; this value is consistent for all four methods within a given block to ensure a fair comparison. The strategies are:
  • Uniform (Matched): Uniformly samples N frames from the video, where N is the output count of AFP for that instance.
  • VSLS (Top-N, Matched): Truncates the VSLS-ranked list to its top N frames.
  • AFP only (Ours): Our pruning method without the semantic graph.
  • AFP + Graph (Full Method, Ours): Our complete method.
Analysis. The results presented in Table 2 reveal a clear and consistent narrative that validates our two-pronged approach. Our full method, AFP + Graph, emerges as the unambiguous winner, achieving the highest accuracy in nearly every tested scenario among all low-budget methods. A deeper analysis reveals two key insights:
First, our clustering-based pruning is inherently superior to naive truncation, even without semantic compensation. This is evident when comparing ‘AFP only‘ with ‘VSLS (Top-N, Matched)‘. Despite operating on an identical frame and token budget, ‘AFP only‘ frequently outperforms ‘VSLS (Top-N, Matched)‘ (e.g., with Qwen2.5-VL-7B-Instruct on VideoMME, and GPT-4o on VideoMME). This demonstrates that our AFP algorithm selects a more diverse and representative set of frames, proving that intelligently eliminating redundancy is more effective than simply selecting the top-ranked frames.
Second, the semantic graph provides a massive performance boost for a negligible cost increase. The consistent and significant performance leap from ‘AFP only‘ to ‘AFP + Graph ‘ across all models and datasets underscores the crucial role of our semantic graph. For instance, with GPT-4o on VideoMME, the addition of the graph elevates the accuracy on short videos from 57.7% to 63.4% (+5.7 points), while only adding a few dozen tokens. This demonstrates that the cost increase is not the primary driver of the performance gain; rather, the gain comes from the highly effective semantic compensation. As shown in our qualitative analysis (Figure 8), the pruned frames preserve the key visual narrative, while the semantic graph provides the high-level context needed to ground the final reasoning.
In conclusion, our two components are synergistically combined: AFP provides a high-quality, non-redundant visual foundation, and the lightweight textual semantic graph provides the crucial, low-cost semantic boost to unlock its full potential. This confirms that the superior performance of our full method stems from its intelligent design, not merely a minor increase in cost.
Figure 8: A qualitative example from a dynamic scene illustrating the effectiveness of Adaptive Frame-Pruning (AFP). For the complex query ”which object is moving quickly?”, the baseline VSLS method selects 16 frames with significant redundancy (e.g., 12 nearly identical initial frames). Our AFP method consolidates these into 3 distinct, representative frames that preserve the key motion cues. This compact set, augmented by the concise semantic graph, provides an efficient yet comprehensive context for the MLLM to derive the correct answer, ”Ambulance”.
Qualitative Analysis
To provide a more intuitive understanding of our method’s ability to handle complex, dynamic scenes beyond quantitative metrics, Figure 8 presents a challenging qualitative example. The task is to identify a fast-moving object based on a highly descriptive and distractor-rich query involving multiple vehicles, crowds, and lights. The baseline VSLS method, aiming for comprehensive coverage, selects 16 keyframes to address this query. However, as shown in the figure, these frames are plagued by severe redundancy that exemplifies the ‘visual echoes’ problem: the first twelve frames are nearly identical, capturing a static, uninformative view of an ambulance. The subsequent three frames show only a slight shift in perspective but remain largely similar, while only the final frame introduces a new contextual element (a tent). This flood of repetitive information risks overwhelming the MLLM and diluting the truly critical visual cues.
Our AFP algorithm excels in such scenarios. It correctly identifies and merges the two large groups of ‘visual echoes’—the 12 static shots and the 3 slightly shifted shots—into one representative frame for each. The final, visually distinct frame is preserved as its own cluster. Consequently, the input is drastically reduced from 16 frames to a highly efficient set of just 3. These three frames, by capturing the ambulance’s initial state, a subtle movement, and the surrounding context, preserve the critical temporal cues needed to infer motion. When this compact visual input is augmented with the generated textual semantic graph, which explicitly highlights the ambulance and its attributes, the MLLM is provided with a sufficient, non-redundant context to correctly identify the ”Ambulance” as the fast-moving object. This example visually demonstrates how AFP distills the essential narrative from a noisy, redundant input.
Conclusion
In this work, we addressed the critical challenge of high token consumption in MLLM-based Video-QA, motivated by the ”less is more” principle. We introduced a novel, two-pronged framework of Adaptive Frame-Pruning (AFP) and text-based semantic graph integration. Our extensive experiments across multiple MLLMs demonstrated that this dual strategy not only drastically reduces frame and token counts but also often improves downstream QA accuracy by mitigating context dilution, particularly for open-source models.
Despite the promising results, we acknowledge limitations. The efficacy of our AFP algorithm depends on visual redundancy in the initial keyframe set; its benefit diminishes in highly dynamic, non-repetitive content. Future work could explore end-to-end joint optimization of selection and pruning, or more advanced MLLM-based graph generation. Ultimately, the principles of our framework could extend to other high-cost video understanding tasks, paving the way for more scalable and accessible video AI.
Supplementary Material
Appendix A A. Complete Ablation Study Results
This section provides the complete and unabridged results of our component-wise ablation study, fulfilling the promise made in the main paper. These comprehensive results demonstrate the robustness of our findings across all models, datasets, and initial keyframe settings. The methods compared are:
  • Baselines: The two high-cost methods using a fixed number of frames (32 or 8), either via ‘Uniform Sampling‘ or ‘VSLS ‘.
  • Matched-Budget Strategies: Four low-cost methods that all operate on the same, drastically reduced number of frames determined by our AFP algorithm for each Video-QA instance.
    • Uniform (Matched): Naive uniform sampling.
    • VSLS (Top-N, Matched): A strong baseline that truncates the VSLS list.
    • AFP only (Ours): Our pruning method without the graph.
    • AFP + Graph (Ours): Our full proposed method.
The detailed results are presented in Table 3 and Table 4. The data consistently shows that under a matched, low-frame budget, our full method (‘AFP + Graph ‘) almost always achieves the best performance among the compressed strategies, validating the synergistic effect of our two core contributions.
Dataset Model Method Avg. Frames Long (%) Medium (%) Short (%) Evaluation starting from 32 Keyframes LongVideoBench GPT-4o Baseline (Uniform) 32.0 53.8 56.5 74.0 Baseline (VSLS) 32.0 54.2 60.0 76.0 Uniform (Matched) 4.2 47.9 48.8 64.0 VSLS (Top-N, Matched) 4.2 50.9 53.8 64.0 AFP only (Ours) 4.2 45.5 47.7 66.0 AFP + Graph (Ours) 4.2 49.4 51.5 80.0 LongVideoBench Qwen2.5-VL-7B Baseline (Uniform) 32.0 32.7 36.5 50.0 Baseline (VSLS) 32.0 38.7 42.3 54.0 Uniform (Matched) 4.2 45.8 45.0 68.0 VSLS (Top-N, Matched) 4.2 50.3 53.1 66.0 AFP only (Ours) 4.2 40.2 41.9 60.0 AFP + Graph (Ours) 4.2 42.9 46.9 66.0 LongVideoBench LLaVA-Video-7B Baseline (Uniform) 32.0 42.3 45.8 54.0 Baseline (VSLS) 32.0 41.7 48.1 54.0 Uniform (Matched) 4.2 38.1 47.3 52.0 VSLS (Top-N, Matched) 4.2 40.2 45.8 52.0 AFP only (Ours) 4.2 41.7 43.5 50.0 AFP + Graph (Ours) 4.2 45.2 50.0 62.0 Evaluation starting from 8 Keyframes LongVideoBench GPT-4o Baseline (Uniform) 8.0 47.1 49.4 67.3 Baseline (VSLS) 8.0 51.2 58.9 74.0 Uniform (Matched) 2.2 47.6 48.1 62.0 VSLS (Top-N, Matched) 2.2 50.3 48.1 66.0 AFP only (Ours) 2.2 45.5 51.5 56.0 AFP + Graph (Ours) 2.2 47.3 53.5 84.0 LongVideoBench Qwen2.5-VL-7B Baseline (Uniform) 8.0 41.1 43.1 62.0 Baseline (VSLS) 8.0 45.8 49.2 54.0 Uniform (Matched) 2.2 38.1 41.9 70.0 VSLS (Top-N, Matched) 2.2 44.0 47.3 56.0 AFP only (Ours) 2.2 40.2 45.0 68.0 AFP + Graph (Ours) 2.2 42.6 45.4 62.0 LongVideoBench LLaVA-Video-7B Baseline (Uniform) 8.0 42.0 46.5 50.0 Baseline (VSLS) 8.0 42.3 46.9 50.0 Uniform (Matched) 2.2 39.3 46.2 42.0 VSLS (Top-N, Matched) 2.2 40.5 42.7 46.0 AFP only (Ours) 2.2 41.1 44.6 44.0 AFP + Graph (Ours) 2.2 45.5 49.6 72.0
Table 3: Complete component-wise ablation study results on the LongVideoBench dataset. For each model and initial frame setting, we compare our full method against various baselines. Bold indicates the best performance among the four matched-budget strategies for each accuracy column.
Dataset Model Method Avg. Frames Long (%) Medium (%) Short (%) Evaluation starting from 32 Keyframes VideoMME GPT-4o Baseline (Uniform) 32.0 55.2 61.0 71.4 Baseline (VSLS) 32.0 55.2 61.9 71.9 Uniform (Matched) 4.3 52.3 52.6 57.1 VSLS (Top-N, Matched) 4.3 49.7 50.2 54.1 AFP only (Ours) 4.3 52.5 52.5 57.7 AFP + Graph (Ours) 4.3 53.0 56.4 63.4 VideoMME Qwen2.5-VL-7B Baseline (Uniform) 32.0 37.5 39.9 54.1 Baseline (VSLS) 32.0 37.9 50.0 55.8 Uniform (Matched) 4.3 39.1 43.5 43.6 VSLS (Top-N, Matched) 4.3 37.9 40.2 42.5 AFP only (Ours) 4.3 38.3 39.8 42.5 AFP + Graph (Ours) 4.3 39.9 44.0 51.8 VideoMME LLaVA-Video-7B Baseline (Uniform) 32.0 35.9 36.5 37.4 Baseline (VSLS) 32.0 36.9 39.0 39.5 Uniform (Matched) 4.3 39.6 40.6 41.7 VSLS (Top-N, Matched) 4.3 38.1 41.1 41.2 AFP only (Ours) 4.3 39.8 40.9 41.5 AFP + Graph (Ours) 4.3 43.8 47.9 56.7 Evaluation starting from 8 Keyframes VideoMME GPT-4o Baseline (Uniform) 8.0 55.2 60.2 69.6 Baseline (VSLS) 8.0 56.9 60.7 68.2 Uniform (Matched) 2.1 52.4 53.4 58.6 VSLS (Top-N, Matched) 2.1 50.5 49.8 56.0 AFP only (Ours) 2.1 53.1 51.8 57.3 AFP + Graph (Ours) 2.1 53.5 56.5 63.8 VideoMME Qwen2.5-VL-7B Baseline (Uniform) 8.0 38.0 47.3 55.4 Baseline (VSLS) 8.0 43.2 49.6 60.8 Uniform (Matched) 2.1 38.5 40.2 42.9 VSLS (Top-N, Matched) 2.1 38.1 37.9 41.4 AFP only (Ours) 2.1 39.2 38.8 42.4 AFP + Graph (Ours) 2.1 38.1 40.2 50.1 VideoMME LLaVA-Video-7B Baseline (Uniform) 8.0 38.0 39.7 38.2 Baseline (VSLS) 8.0 38.5 42.7 38.1 Uniform (Matched) 2.1 40.6 41.2 38.3 VSLS (Top-N, Matched) 2.1 39.2 40.2 38.5 AFP only (Ours) 2.1 38.9 41.0 40.3 AFP + Graph (Ours) 2.1 44.2 47.9 54.5
Table 4: Complete component-wise ablation study results on the VideoMME dataset. Format and highlighting rules are the same as in Table 3.
Appendix B B. Further Implementation Details
This section provides in-depth details about our implementation to ensure full reproducibility of our work.
B.1. AFP Algorithm Details
Our Visual Echo Suppression (AFP) algorithm is implemented in Python using the scikit-learn library (Pedregosa et al. 2011) for clustering. Below are the key implementation details that complement the description in the main paper’s Methodology section.
Clustering Parameters.
For the core clustering step, we use the AgglomerativeClustering class from scikit-learn. The linkage parameter, which determines the merge strategy, is set to ’average’. This means the distance between two clusters is defined as the average of the distances between all pairs of samples, with one sample from each cluster.
Small Cluster Refinement.
To enhance the robustness of our clustering, we implement a refinement step (refine_clusters in our script). After the initial clustering, any cluster containing fewer than two frames is considered unstable. Such singleton clusters are merged into their nearest neighboring cluster, where proximity is determined by the average visual cosine distance between the frames of the two clusters. This prevents trivial clusters and ensures a more meaningful grouping.
Adaptive Threshold Calculation.
As mentioned in the main paper, the distance_threshold for clustering is determined adaptively. Specifically, after calculating all pairwise visual cosine distances, we fit a Gaussian Kernel Density Estimator (KDE) to their distribution. The threshold is then calculated as $\tau=p+0.15$, where $p$ is the distance value corresponding to the peak of the density function. The constant offset of 0.15 is an empirically chosen value designed to prevent the clustering from being overly conservative (i.e., creating too many small clusters) and to encourage the merging of closely related ‘visual echoes’.
B.2. Prompt Structure and MLLM Inference
The structure of the prompts sent to the MLLM is critical for achieving consistent and reproducible results. In this section, we detail the two primary prompt templates used in our framework: one for our versatile semantic graph generation (Case 2), and one for the final downstream Video-QA task.
Prompt for Semantic Graph Generation.
For our low-cost, universally applicable semantic graph generation strategy (Case 2), we employ a sophisticated prompt designed to elicit rich, structured information from the MLLM based solely on textual input. As detailed in Figure 9, the prompt is structured using several advanced prompting techniques:
  • Persona Prompting: We assign the MLLM the role of an ”expert visual scene analyst” to activate its advanced reasoning capabilities.
  • Chain-of-Thought Instructions: The prompt provides a clear, step-by-step thinking process (Key Entity Identification, Contextual Cue Identification, Relationship Extraction) for the MLLM to follow.
  • Open-ended yet Structured Relation Extraction: Crucially, unlike previous methods like VSLS that pre-define a closed set of relation types, our prompt encourages flexibility. It explicitly defines two categories of relations—Concrete (e.g., verbs, prepositions) and Abstract Logical (e.g., ‘spatial‘, ‘causal‘)—and provides clear examples and definitions for each. This ”dual-track” design allows for the capture of a much richer and more nuanced set of semantic connections while maintaining a structured output.
  • Few-Shot Example: A comprehensive example is provided to demonstrate the desired input-output format and the expected quality of the extracted graph.
The full prompt is shown in Figure 9.
Figure 9: The full prompt template for LLM-based semantic graph generation. This prompt leverages persona, chain-of-thought, and open-ended relation extraction to guide the LLM in deconstructing the query.
Prompt for Downstream Video-QA.
For the final question-answering stage, we utilize a clean, direct, and robust prompt template, detailed in Figure 10. This prompt is designed for maximum compatibility across different MLLMs. It clearly presents the visual evidence (via ’<image>’ placeholders), followed by the optional textual semantic graph, and finally the question and multiple-choice options. The instructions are direct and unambiguous, strictly constraining the model to output a single uppercase letter corresponding to its chosen answer. This minimalist design minimizes prompt engineering sensitivity and ensures that performance differences are attributable to the quality of the visual and semantic inputs, rather than variations in prompt interpretation.
Figure 10: The full MLLM prompt template used for the final Video-QA task. It integrates image placeholders, the optional semantic graph, and the QA content into a single, direct query.
B.3. Computational Environment
Our experiments were conducted within the computational environment detailed in Table 5. The deep learning models were implemented using the PyTorch framework (Paszke et al. 2019).
Component Specification
Language Python 3.9+
Deep Learning Framework PyTorch
Key Libraries scikit-learn, transformers, OpenCV
Conda Environment Name VC
GPU 8x NVIDIA RTX A6000 (48GB VRAM each)
CUDA Version 12.1
Table 5: Software and hardware environment for our experiments.
B.4. Representative Frame Selection Strategy
A crucial step in our AFP algorithm is selecting a single representative frame from each generated cluster. To determine the most effective and robust approach, we conducted a comprehensive ablation study comparing three distinct strategies. The experiments were performed on the LongVideoBench dataset with the GPT-4o model, starting from Top-8 keyframes provided by VSLS, and the results reported are averaged over multiple runs to ensure reliability. This analysis focuses on the ‘AFP only‘ setting (without the semantic graph) to purely evaluate the quality of the selected visual information.
  • Score-based (Highest VSLS Score): This strategy leverages external information by selecting the frame from each cluster that has the highest initial relevance score assigned by the upstream VSLS selector.
  • Centroid-based (Visual Centroid): This strategy is self-contained. It selects the ”centroid frame”—the frame with the minimum average feature distance to all other frames within the same cluster, making it the most visually representative.
  • Relevance-based (Highest Query Similarity): This strategy introduces task-specific guidance by selecting the frame with the highest CLIP similarity score to the ‘Question + Options‘ text prompt.
Experimental Settings Averaged Results
Dataset Model Frame Source Selection Strategy Avg. Acc (%) Avg. Frames
LongVideoBench GPT-4o VSLS (Top-8) Score-based 47.14 2.1
Centroid-based 48.53 2.1
Relevance-based 46.67 2.2
Table 6: Ablation study on representative frame selection strategies. The experiment was conducted in the ‘AFP only‘ setting to purely evaluate the quality of the selected visual information. Results are averaged over multiple runs. The Centroid-based strategy demonstrates the best performance.
As shown in Table 6, the averaged results over multiple runs confirm that the Centroid-based strategy holds a clear advantage. It achieves the highest overall accuracy (48.53%), outperforming both the Score-based and Relevance-based approaches.
The superiority of the Centroid strategy is theoretically sound, especially in the absence of a semantic graph. By selecting the frame closest to the cluster’s feature-space center, it guarantees the most visually representative frame is chosen, which is the most robust way to minimize information loss during the pruning process. This self-contained logic is arguably more generalizable than relying on external scores (from VSLS or CLIP), which might be noisy or biased. Based on these comprehensive findings, we adopted the Centroid-based strategy for all main experiments reported in this paper.
Appendix C C. Generalizability Study: Applying AFP to T*
To validate the universal applicability of our AFP framework as a post-processing module, we conducted a comprehensive set of experiments by applying it to the output of another state-of-the-art keyframe selector, T*. This section presents a full generalizability and ablation study across all three tested MLLMs (GPT-4o, Qwen2.5-VL-7B-Instruct, and LLaVA-Video-7B-Qwen2) on both the LongVideoBench and VideoMME datasets.
The methods compared in this section are:
  • T* Baseline: The high-cost method using the original Top-32 or Top-8 frames selected by T*.
  • Matched-Budget Strategies: Four low-cost methods that all operate on the same, drastically reduced number of frames determined by our AFP algorithm for each instance.
    • Uniform (Matched): Naive uniform sampling with a matched frame budget.
    • T* (Top-N, Matched): A strong baseline that truncates the original T* list to match the frame budget.
    • AFP only (Ours): Our pruning method applied to T*’s frames, but without using the semantic graph during QA.
    • AFP + Graph (Ours): Our full proposed method, combining AFP with a generated semantic graph.
A Note on Semantic Graph Generation for T*.
A key point in our experimental design is the source of the semantic graph for the AFP + Graph experiments in this section. While the T* method itself does not explicitly extract semantic-logical relations, our experiments for both T* and VSLS were conducted on the exact same set of Video-QA instances from the LongVideoBench and VideoMME benchmarks. This allows us to directly reuse the semantic graphs that were already generated for the VSLS experiments for each corresponding data instance. This approach is not only highly efficient, maintaining the zero-cost nature of our primary setup, but also ensures a perfectly fair comparison: the semantic graph, as a variable, is held constant, isolating the performance difference to the quality of the initial keyframe sets provided by T* versus VSLS.
The results are presented in Table 7 and Table 8.
Analysis of Generalizability.
The results robustly confirm that our AFP framework is a versatile, plug-and-play refinement module. As shown in the tables, our full method, AFP + Graph, consistently improves performance or offers a superior cost-utility trade-off when applied to the T* baseline. For instance, with GPT-4o on LongVideoBench (from Top-8), our method boosts the accuracy on short videos from 72.7% to 80.0%.
Notably, the performance gains are again most pronounced on the open-source models. Using LLaVA-Video-7B-Qwen2 on VideoMME (from Top-32), our method elevates the accuracy on short videos from 37.7% (T* Baseline) to 55.0%—a staggering 17.3-point absolute improvement—while using only 4.1 frames instead of 32. This reinforces our main finding that open-source MLLMs benefit immensely from our pruning and compensation strategy. Furthermore, the success of the LLM-generated semantic graph in this setting validates its effectiveness as a low-cost, universally applicable alternative when pre-existing semantic information is unavailable. This confirms that the benefits of our framework are not contingent on a specific upstream selector like VSLS, but stem from the fundamental principles of intelligent redundancy reduction and semantic compensation.
Figure 11: Generalizability of AFP + Graph on T* Keyframes. The plot shows the performance-efficiency trade-off on both LongVideoBench (left) and VideoMME (right) datasets across three MLLMs. Solid lines represent the high-cost T* baseline at 8 and 32 frames. Dashed lines represent our low-cost AFP + Graph method applied to the corresponding T* outputs. Our method consistently operates in the top-left (high-accuracy, low-frame-count) region, demonstrating its effectiveness as a universal refinement module.
The results robustly visualized in Figure 11, confirm that our AFP framework is a versatile plug-and-play refinement module. As shown in the plots, our AFP + Graph method (dashed lines) consistently operates in the highly desirable ”top-left” quadrant of the performance-efficiency space. It achieves an accuracy that is either competitive with or, in many cases, significantly superior to the high-cost T* baseline (solid lines), while using only a fraction of the frames. For instance, on LongVideoBench with the LLaVA-Video-7B model, our method applied to T* Top-8 outputs (2.2 frames) achieves a 51.15% accuracy, surpassing the T* Top-8 baseline (48.06%) by over 3 absolute points. This demonstrates that AFP + Graph is not overfitted to the properties of a single selector like VSLS and can effectively enhance other keyframe selection methods.
Dataset Model Method Avg. Frames Long (%) Medium (%) Short (%) Evaluation starting from 32 Keyframes (Source: T* Top-32) LongVideoBench GPT-4o T* Baseline 32.0 53.1 59.4 74.3 Uniform (Matched) 4.2 46.4 48.8 62.0 T* (Top-N, Matched) 4.2 49.1 55.0 62.0 AFP only (Ours) 4.2 45.8 48.8 70.0 AFP + Graph (Ours) 4.2 48.5 53.1 80.0 LongVideoBench Qwen2.5-VL-7B T* Baseline 32.0 38.7 41.9 40.0 Uniform (Matched) 4.2 38.1 43.5 64.0 T* (Top-N, Matched) 4.2 41.4 48.5 68.0 AFP only (Ours) 4.2 41.1 41.2 62.0 AFP + Graph (Ours) 4.2 42.6 43.1 68.0 LongVideoBench LLaVA-Video-7B T* Baseline 32.0 40.2 44.2 50.0 Uniform (Matched) 4.2 38.7 46.9 48.0 T* (Top-N, Matched) 4.2 42.3 46.9 48.0 AFP only (Ours) 4.2 41.1 49.2 44.0 AFP + Graph (Ours) 4.2 42.0 52.7 62.0 Evaluation starting from 8 Keyframes (Source: T* Top-8) LongVideoBench GPT-4o T* Baseline 8.0 51.9 52.4 72.7 Uniform (Matched) 2.2 45.5 48.8 64.0 T* (Top-N, Matched) 2.2 46.4 53.1 64.0 AFP only (Ours) 2.2 47.3 48.8 68.0 AFP + Graph (Ours) 2.2 46.4 52.7 80.0 LongVideoBench Qwen2.5-VL-7B T* Baseline 8.0 42.0 47.7 54.0 Uniform (Matched) 2.2 35.7 38.5 70.0 T* (Top-N, Matched) 2.2 43.8 48.1 56.0 AFP only (Ours) 2.2 38.7 41.9 66.0 AFP + Graph (Ours) 2.2 42.0 45.4 66.0 LongVideoBench LLaVA-Video-7B T* Baseline 8.0 39.6 50.0 48.0 Uniform (Matched) 2.2 40.8 46.5 48.0 T* (Top-N, Matched) 2.2 39.3 45.4 42.0 AFP only (Ours) 2.2 39.6 49.2 52.0 AFP + Graph (Ours) 2.2 45.5 51.5 62.0
Table 7: Generalizability and ablation results of applying our method to the output of the T* keyframe selector on the LongVideoBench dataset. Bold indicates the best performance among the four matched-budget strategies for each accuracy column.
Dataset Model Method Avg. Frames Long (%) Medium (%) Short (%) Evaluation starting from 32 Keyframes (Source: T* Top-32) VideoMME GPT-4o T* Baseline 32.0 59.3 63.5 69.5 Uniform (Matched) 4.1 52.4 55.2 57.7 T* (Top-N, Matched) 4.1 49.8 49.4 54.6 AFP only (Ours) 4.1 51.8 53.6 55.7 AFP + Graph (Ours) 4.1 52.6 55.3 63.1 VideoMME Qwen2.5-VL-7B T* Baseline 32.0 34.9 45.6 55.2 Uniform (Matched) 4.1 39.9 43.1 43.5 T* (Top-N, Matched) 4.1 37.9 40.1 43.4 AFP only (Ours) 4.1 39.6 40.4 44.6 AFP + Graph (Ours) 4.1 39.4 40.7 52.3 VideoMME LLaVA-Video-7B T* Baseline 32.0 35.8 39.6 37.7 Uniform (Matched) 4.1 38.7 42.9 41.3 T* (Top-N, Matched) 4.1 40.2 39.3 40.3 AFP only (Ours) 4.1 38.4 41.7 39.3 AFP + Graph (Ours) 4.1 41.7 47.1 55.0 Evaluation starting from 8 Keyframes (Source: T* Top-8) VideoMME GPT-4o T* Baseline 8.0 55.9 57.3 56.4 Uniform (Matched) 2.1 50.1 54.2 58.6 T* (Top-N, Matched) 2.1 49.9 51.7 56.2 AFP only (Ours) 2.1 52.2 51.1 56.7 AFP + Graph (Ours) 2.1 52.3 55.3 66.1 VideoMME Qwen2.5-VL-7B T* Baseline 8.0 40.3 50.0 54.9 Uniform (Matched) 2.1 37.7 39.4 42.5 T* (Top-N, Matched) 2.1 36.1 38.5 41.4 AFP only (Ours) 2.1 38.0 39.3 41.5 AFP + Graph (Ours) 2.1 36.2 40.4 49.5 VideoMME LLaVA-Video-7B T* Baseline 8.0 37.5 40.4 37.2 Uniform (Matched) 2.1 38.8 41.6 39.2 T* (Top-N, Matched) 2.1 39.3 41.0 38.8 AFP only (Ours) 2.1 37.8 40.4 40.0 AFP + Graph (Ours) 2.1 42.3 46.2 56.1
Table 8: Generalizability and ablation results of applying our method to the output of the T* keyframe selector on the VideoMME dataset. Format and highlighting rules are the same as in Table 7.
Appendix D D. Ablation Study on Prompt Formulation
The formulation of the prompt, particularly how the semantic graph is textualized and how instructions are given to the MLLM, can significantly impact performance. To determine the optimal prompt structure for our main experiments, we conducted an ablation study on several variants. All experiments were performed on the LongVideoBench dataset with the GPT-4o model, starting from Top-8 keyframes.
D.1. Prompt Component Variants
We tested combinations of two graph_context formats and two system_prompt formats:
Graph Context Formats.
  • Concise Triplet (G1): Our final choice. This format is simple and structured, directly presenting nodes and raw triplet relationships (e.g., (object1, relation, object2)).
  • Verbose Natural Language (G2/G3): These formats attempt to convert the graph into more human-like sentences (e.g., “object1 appears with object2”).
System Prompt Formats.
  • Direct Prompt (P1): Our final choice. This is a concise, direct instruction for the QA task.
  • Instructional Prompt (P2): This is a more verbose prompt that assigns an ”expert” persona to the MLLM and provides detailed step-by-step guidelines.
D.2. Results and Analysis
The results of our prompt ablation are presented in Table 9. Our analysis led us to select the combination of Concise Triplet graph (G1) and Direct Prompt (P1) for all main experiments, for the following key reasons:
Graph Prompt Avg. Accuracy (%)
Format Format Long Medium Short
G1 (Ours) P1 (Ours) 47.0 52.5 69.0
G2 P1 46.1 51.7 66.4
G1 P2 46.7 52.9 70.7
G2 P2 45.9 49.9 69.0
G3 P2 44.9 51.2 70.7
Table 9: Ablation study on prompt formulation. Results are averaged over multiple runs. Our chosen combination (G1, P1) is highlighted in bold.
  • Robust and Stable Performance: While other combinations, such as (G1, P2), achieve marginally higher accuracy on certain video lengths, the (G1, P1) combination demonstrated the most stable and consistent high performance across all our preliminary and main experiments. As noted in our experimental logs, other verbose combinations frequently yielded unexpectedly low results in some runs, indicating a lack of robustness. The (G1, P1) setting, in contrast, reliably produced strong results.
  • Token Efficiency: The G1 and P1 formats are significantly more concise than their verbose counterparts. This results in a lower token count for each query, which directly aligns with our paper’s core objective of maximizing token efficiency.
  • Simplicity and Generalizability: The direct, structured format of (G1, P1) provides clear instructions to the MLLM without excessive ”prompt engineering.” We believe this simpler format is more likely to generalize well across different MLLMs, as it relies on fundamental instruction-following capabilities rather than sensitivity to nuanced persona-based instructions.
In conclusion, our chosen (G1, P1) prompt formulation represents the best trade-off between performance, stability, and token efficiency, making it the most suitable choice for our proposed method.
Appendix E E. Analysis of Hyperparameters
This section provides the detailed data and analysis supporting our hyperparameter selection, complementing the cost-utility trade-off visualization presented in the main paper. We conducted a sensitivity analysis on the two key parameters of our AFP algorithm: the feature fusion ratio $\alpha$ and the distance metric weight $\beta$. All experiments were performed on the LongVideoBench dataset using the Qwen2.5-VL-7B-Instruct model, starting from Top-8 keyframes. We report the averaged results over multiple runs in Table 10.
Hyperparameter Setting Avg. Accuracy (%) Avg. Frames
Analysis of Feature Fusion Ratio $\alpha$ (fixing $\beta=0.9$)
$\alpha$ = 0.2 40.02 2.31
$\alpha$ = 0.4 39.83 2.31
$\alpha$ = 0.6 (Ours) 41.58 2.80
$\alpha$ = 0.8 41.04 2.31
Analysis of Distance Metric Weight $\beta$ (fixing $\alpha=0.6$)
$\beta$ = 0.3 41.00 3.10
$\beta$ = 0.5 40.30 3.00
$\beta$ = 0.7 41.00 3.00
$\beta$ = 0.9 (Ours) 41.58 2.80
Table 10: Detailed sensitivity analysis of hyperparameters $\alpha$ and $\beta$. Results are averaged over multiple runs on the LongVideoBench dataset using the Qwen2.5-VL-7B-Instruct model (from Top-8). Our chosen parameters ($\alpha=0.6,\beta=0.9$‘) are highlighted in bold.
Analysis and Selection Strategy.
Our analysis reveals a clear cost-utility trade-off. For the feature fusion ratio $\alpha$, the results show that $\alpha=0.6$ achieves the highest average accuracy of 41.58%, suggesting an optimal balance between visual and semantic features. For the distance metric weight $\beta$, a higher value (e.g., 0.9) leads to more aggressive clustering (fewer frames) and, in this case, also the highest accuracy.
This demonstrates that the hyperparameters of AFP are not arbitrary fixed values, but rather tunable ”knobs” that control the cost-utility balance. Based on this systematic analysis, we selected the configuration of $\alpha=0.6$ and $\beta=0.9$ for all main experiments, as it consistently provided the best overall performance and a favorable efficiency-accuracy balance in our development set.
Appendix F D. Additional Qualitative Examples
To demonstrate the versatility and robustness of our Adaptive Frame-Pruning (AFP) method beyond the case presented in the main paper, this section provides additional qualitative examples from different types of Video-QA tasks from the VideoMME dataset. These examples illustrate how AFP effectively handles various forms of visual redundancy while preserving the crucial information needed to answer diverse queries.
Figure 12: Qualitative example for a counting task (Video ID: N1cdUjctpG8). The query asks for the number of porcelain jars. The baseline VSLS method selects 32 frames, many of which are redundant, capturing similar views of the jars. Our AFP algorithm prunes this set to just 4 frames by consolidating repetitive views, enabling the MLLM to perform an accurate count with a highly efficient prompt.
Example 1: Counting Task.
Figure 12 showcases a counting-based question: “How many porcelain jars were discovered…?”. To answer this, the model must identify and aggregate all visible jars. The baseline VSLS selector, in its effort to be comprehensive, retrieves 32 keyframes. Among these, at least seven frames capture the porcelain jars from slightly different but highly overlapping angles, creating significant redundancy. Our AFP algorithm correctly identifies these frames as a single ‘visual echo’. It merges these seven redundant views into a single, clear representative frame that encapsulates the essential information for counting. By combining this consolidated view with three other contextually important (but non-redundant) frames, AFP reduces the input from 32 to just 4 frames. This compact set is sufficient for the MLLM to correctly answer ”9”, demonstrating how our method eliminates noise while preserving the core evidence required for aggregation tasks.
Figure 13: Qualitative example for a fine-grained detail recognition task (Video ID: XWfqBKeC0g8). The query requires identifying small text within the frame. VSLS selects 30 frames, with the first six being nearly identical. Our AFP intelligently prunes this to 3 frames, critically retaining two distinct frames that both contain the ”2000x” text, ensuring this vital but subtle information is preserved for the MLLM.
Example 2: Fine-Grained Detail Recognition.
Figure 13 illustrates a more challenging task requiring the recognition of fine-grained details—specifically, small text in the query “What level of magnification is used…?”. The crucial text ”2000x magnification” appears in multiple early frames. The VSLS baseline selects 30 frames, with the first six being almost perfect duplicates. A naive de-duplication might aggressively merge all six into one. However, our AFP algorithm demonstrates a more nuanced approach. It prunes the 30 initial frames down to just 3. Critically, among these three, it preserves two of the initial six frames. This is because subtle changes in focus or cellular movement create enough feature variance for our adaptive clustering to treat them as distinct (though closely related) pieces of evidence. This intelligent, slightly conservative pruning proves highly effective: it drastically reduces redundancy while increasing the robustness of the final answer by providing the MLLM with multiple opportunities to spot the critical, fine-grained text. This case highlights that AFP does not merely de-duplicate, but rather optimizes the trade-off between redundancy reduction and evidence preservation.
Appendix G G. Discussion of Limitations and Failure Cases
To provide a balanced perspective on our work, this section honestly discusses the inherent limitations of our proposed method and analyzes two representative failure cases. We believe this discussion is crucial for understanding the boundaries of our approach and for inspiring future research in efficient video understanding.
G.1. General Limitations
Our approach, while effective, is subject to two limitations:
  • Dependency on Redundancy. The core premise of Adaptive Frame-Pruning (AFP) is the existence of visual redundancy in the initial keyframe set. In scenarios featuring highly dynamic, non-repetitive content, such as a montage of rapidly changing scenes, the opportunities for clustering are naturally diminished. In such cases, AFP would correctly identify most frames as unique and perform minimal reduction, thus offering limited efficiency gains.
  • Ceiling Effect from Upstream Selector. Our method operates as a post-processing refinement layer. Consequently, its performance is fundamentally capped by the quality of the initial keyframe set provided by the upstream selector (e.g., VSLS). If the initial selector fails to retrieve the frames containing the necessary evidence to answer a question, AFP cannot recover this information, as its role is to prune, not to discover.
G.2. Failure Case Analysis
Beyond general limitations, we present two failure cases that highlight specific challenges.
Case 1: Loss of Fine-Grained Information during Pruning.
The first case, shown in Figure 14, demonstrates a scenario where our AFP algorithm, while correctly identifying visual similarity, inadvertently discards critical, fine-grained information. The question asks which of human’s ancestors first began to walk on two legs. The initial 32 frames selected by VSLS correctly contain frames depicting both ”Ramapithecus” (the correct answer, appearing earlier) and ”Ardipithecus Ramidus”. However, because the overall visual composition of these frames—the posture of the ancestor, the background—is highly similar, our algorithm, which relies on global ResNet and CLIP features, perceives them as a single ‘visual echo’. Consequently, it merges them into one cluster and, in this instance, retains only the frame for ”Ardipithecus Ramidus”. The MLLM, presented with this incomplete evidence, understandably arrives at the wrong conclusion. This failure highlights a key limitation: our global feature-based approach can be insensitive to subtle but semantically decisive local details, such as small pieces of text. This points towards a clear avenue for future improvement: integrating more localized feature extractors, such as Optical Character Recognition (OCR) models, into the clustering process.
Figure 14: An analysis of a failure case caused by information loss during pruning. The initial keyframes from VSLS contain the evidence for the correct answer ”Ramapithecus” (highlighted in green). However, due to high visual similarity with other frames, our AFP algorithm incorrectly prunes this crucial frame, leading the VLLM to a wrong answer based on the remaining evidence.
Figure 15: An analysis of a failure case caused by the insufficiency of the upstream selector. The question requires a global understanding of a narrative. The initial 32 frames from VSLS are already insufficient to answer correctly. Our AFP method, while efficiently pruning the input, cannot recover the missing information, leading to an inevitable failure by the VLLM. This illustrates the performance ceiling imposed by the initial keyframe selection.
Case 2: Insufficiency of the Upstream Selector.
The second case, shown in Figure 15, exemplifies the ”Ceiling Effect” limitation. The question asks for the main reason people commemorate Qu Yuan, a query that requires a deep, holistic understanding of a narrative spread across the entire video, likely conveyed through multiple subtitles. The initial 32 frames selected by VSLS, while topically relevant, fail to capture the full narrative context required to make such a nuanced judgment. As a result, even before our AFP algorithm is applied, the provided visual information is insufficient for the MLLM to correctly answer the question. After our method prunes the already insufficient set of frames, the MLLM still fails. This case clearly demonstrates that our method’s success is contingent on the upstream selector providing a set of frames that, even if redundant, collectively contains the necessary information. It highlights a broader challenge for all sparse sampling methods when dealing with questions that require global, narrative-level reasoning.
Summary and Future Directions.
These failure cases point towards a clear avenue for future improvement. The primary weakness identified is the reliance on global features, which can overlook fine-grained, decisive details like small text or subtle object differences. A promising future direction would be to develop a more hierarchical or adaptive feature fusion mechanism. For instance, the algorithm could first perform a global clustering as it does now. Then, for clusters with high internal visual similarity but containing potentially important textual cues (identified via a lightweight OCR scan), it could dynamically switch to a more localized, patch-based feature representation to re-evaluate the clustering decision. This would allow the framework to be both efficient at a global level and precise at a local level when necessary, enhancing its robustness and accuracy on a wider range of complex queries.